OpenStudy (anonymous):

(I just had my surgery and I can't focus well, meds are kicking in, help me. ) An energy bar company manufactures peanut butter and chocolate bars. Peanut butter bars cost 25¢ per lb. The chocolate bars cost 30¢ per lb. The company sold 260 lbs of bars for $71. How many of the bars sold were peanut butter?

OpenStudy (anonymous):

284

OpenStudy (anonymous):

Explain please

OpenStudy (anonymous):

k I divided it 71/0.25

OpenStudy (anonymous):

That doesn't help me at all..

OpenStudy (anonymous):

Let peanut butter bars sold be x lb cost of it 25¢ per lb. = 25x Then chocolate bars sold = (260 – x) cost 30¢ per lb = 30(260 – x) total cost = 25x + 30(260 – x) = 7100 25x + 7800 – 30x = 71 5x = 700 x = 140 -------- peanut butter bars 260 – 140 = 120 chocolate bars

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

need anything else?

OpenStudy (anonymous):

no thank you.

OpenStudy (anonymous):

okay haha. and I hope u get better soon!